The answer is suppose to be: 6.67m/s

Could you break it down step by step, for I keep getting an answer of 9

100 = [-10242V4 - 2002V, ] +2vi
Answer #1

lo0= (-102+22- (2X)0X2V2)) +2k %3D floo +25- fol2 +2 4) 12 2v- 38V2 -200=0 V 19v2-100 =0 Vz = -19± 19+(2x IX100) 2. và= 23.